OpenStudy (anonymous):

evaluate this function: w(t)=3t+2; Find w(2-t)

jimthompson5910 (jim_thompson5910):

w(2-t) means "wherever you see a 't' replace it with '2-t' and simplify"

jimthompson5910 (jim_thompson5910):

w(t)=3t+2 w(2-t)=3(2-t)+2 I'll let you simplify

jimthompson5910 (jim_thompson5910):

Tell me what you get when you do

OpenStudy (anonymous):

okay good, i am just making sure my answer is correct. i got w(2-t)=8-3t

jimthompson5910 (jim_thompson5910):

you got it, the answer is w(2-t) = 8-3t or w(2-t) = -3t+8

OpenStudy (anonymous):

i have another one that i need help with. k(t)=\[\left| x+3 \right|\] Find k(-4t)

jimthompson5910 (jim_thompson5910):

hmm did you mean \[\Large k(t) = \left| t+3 \right|\] Notice how it's t+3 and NOT x+3

OpenStudy (anonymous):

oh i am so sorry. that isn't even right. i accidently combined two problems when i glanced at my worksheet. it is supposed to be k(t)= t+4; Find k(-4t)

jimthompson5910 (jim_thompson5910):

oh ok, that makes more sense

jimthompson5910 (jim_thompson5910):

k(t)= t+4 k(-4t)= (-4t)+4 k(-4t)= -4t+4

jimthompson5910 (jim_thompson5910):

That's all there is to it really

OpenStudy (anonymous):

that is what i got, but it seemed too easy like i was forgetting something. thanks! :)

jimthompson5910 (jim_thompson5910):

yw